# EmoFace: Emotion-Content Disentangled Speech-Driven 3D Talking Face with Mesh Attention

### v1

@@ Abstract
The creation of increasingly vivid 3D virtual digital humans has become a hot topic in recent years. Currently, most speech-driven work focuses on training models to learn the relationship between phonemes and visemes to achieve more realistic lips. However, they fail to capture the correlations between emotions and facial expressions effectively. To solve this problem, we propose a new model, termed EmoFace. EmoFace employs a novel Mesh Attention mechanism, which helps to learn potential feature dependencies between mesh vertices in time and space. We also adopt, for the first time to our knowledge, an effective self-growing training scheme that combines teacher-forcing and scheduled sampling in a 3D face animation task. Additionally, since EmoFace is an autoregressive model, there is no requirement that the first frame of the training data must be a silent frame, which greatly reduces the data limitations and contributes to solve the current dilemma of insufficient datasets. Comprehensive quantitative and qualitative evaluations on our proposed high-quality reconstructed 3D emotional facial animation dataset, 3D-RAVDESS ($5.0343\times 10^{-5}$mm for LVE and $1.0196\times 10^{-5}$mm for EVE), and publicly available dataset VOCASET ($2.8669\times 10^{-5}$mm for LVE and $0.4664\times 10^{-5}$mm for EVE), demonstrate that our algorithm achieves state-of-the-art performance.
Introduction
Generating realistic 2D/3D face animations has received great attention in a number of fields, including film production, computer games, virtual reality, education, etc. due to the great importance of sound and facial expressions in the communication of human beings. As a rule-based approach (Edwards et al. 2016; Taylor et al. 2012; Xu et al. 2013), the successful application of JALI (Edwards et al. 2016) in game production signifies the potential commercial value of 3D face animation generation. Recently, several 3D facial animation models based on deep learning (Karras et al. 2017; Cudeiro et al. 2019; Richard et al. 2021; Fan et al. 2022; Xing et al. 2023; Peng et al. 2023a; Wu et al. 2023; Stan, Haque, and Yumak 2023) have been possible to produce impressive 3D face animations. Moreover, these models can also significantly save time and labour costs compared to manual production, thus being preferred for academic research and commercial exploration.
Figure 1: Input the speech with different emotions, EmoFace will generate realistic 3D talking faces. Audio corresponds to the mesh in the same position, e.g. audio with the emotion ”Angry” corresponds to the upper left corner of the mesh.
However, there are still some situations that have not been adequately addressed, for example, the relationship between emotions and facial expressions. Researches on 2D talking face generation (Ji et al. 2021; Sinha et al. 2022) have focused on this aspect, while there is still scarce on emotional speech-driven 3D facial animation. Most of the preceding speech-driven methods focus more on achieving high-quality lip synchronisation, ignoring the impact of emotion on facial motions. MeshTalk (Richard et al. 2021) and the method proposed by Wu et al. (2023) note the importance of facial expressions for the overall full-face synthesis effect, but do not include emotion as a concern. EmoTalk (Peng et al. 2023b) and EMOTE (Daněček et al. 2023) recognize the significance of emotion, thus disentangling emotional and content information in speech to generate higher-quality emotional 3D facial animations. EmoTalk disentangles content information but does not use it as a driving source for the transformer decoder module, and the training process must start from a silent state, limiting the scope of the dataset. EMOTE is a two-stage model based on FLAME (Li et al. 2017) parameters rather than mesh vertices and does not achieve end-to-end training. Theoretically, vertex-based methods can more accurately control the motion of each mesh vertex than the parameter-based methods since coefficient estimation is a highly abstract problem that ignores spatial correlations (Guan Ming, Prayook, and Wei Tech 2020).
@@ Introduction
To address the above problems, in this paper, we propose a novel end-to-end emotional speech-driven 3D face animation approach, as shown in Figure 1. We first disentangle emotion and content from speech and extract their features separately, then a latent space decoder is used to obtain the emotion-based and content-based vertex offsets, and eventually we integrate them using Mesh Attention to predict the final offsets. Due to that the fusion weights for the emotion branch and the content branch are variable in both temporal and spatial domains, we add a 3D spiral convolution operator in Mesh Attention for effective feature exaction. Moreover, during the training process, it is observed that the autoregression scheme of the transformer decoder results in error accumulation, which ultimately leads to a deterioration in the output and increases the training time. Although, existing teacher-forcing scheme can help to alleviate this problem, exposure bias may occur when switching from training to inference since no access to real historical data. Inspired by scheduled sampling (Mihaylova and Martins 2019), we propose a self-growing scheme that gradually reduces the ratio of groundtruth provided during training, transitioning from teacher-forcing to scheduled sampling. In order to train our emotion-content disentanglement model better, we use SOTA method EMOCA (Daněček, Black, and Bolkart 2022) to reconstruct reliable 3D faces from 2D dataset RAVDESS (Livingstone and Russo 2018). Meanwhile, we employ Kalman Filter (Kalman 1960) to ensure smoother vertex motions and improve the quality of reconstruction. Extensive qualitative and quantitative experiments demonstrate that our method outperforms current state-of-the-art methods for better generation of facial expressions and lip synchronisation.
In summary, our main contributions are as follows:
@@
  • We provide a new high-quality emotional speech-driven dataset, 3D-RAVDESS, and a pipeline for acquiring 4D datasets.
  • We introduce emotion and content feature disentangling encoder and intermediate supervision to predict offset gains from long-term emotion and short-term content features, respectively.
@@
  • We present a novel Mesh Attention with SpiralConv3D that efficiently extracts features in the temporal-spatial domain, dynamically fusing the contributions of emotion and content branches to the offset of vertices.
  • We propose a self-growing training scheme combining teacher-forcing and scheduled sampling to gradually increase the difficulty of prediction and improve the robustness of the model. Extensive experiments and the user study demonstrate the superiority of EmoFace over existing SOTA methods in terms of realistic facial animation, emotion expression and lip synchronisation on both 3D-RAVDESS and VOCASET.
Related Work
Speech-Driven 3D Talking Face
@@
Speech-driven 3D talking face generation is a task to generate realistic facial animations based on speech (Cao et al. 2005; Karras et al. 2017; Lahiri et al. 2021; Pham, Cheung, and Pavlovic 2017; Taylor et al. 2017; Zhang et al. 2010). A number of rule-based approaches have been employed, including the capture of the relationship between visemes (Mattheyses and Verhelst 2015) and facial action units (FAUs) (Ekman and Friesen 1978), as well as the establishment of a phoneme-viseme mapping that has yielded highly promising results in early studies (Edwards et al. 2016; Taylor et al. 2012; Xu et al. 2013; Massaro et al. 2001). Nevertheless, these methods do not easily transfer to new faces, heavily rely on manual crafting, and consume an excessive amount of time.
Figure 2: The overall framework of EmoFace. The emotion and content branches disentangle the information in speech, while Mesh Attention fuses these two branches to obtain the final result. The entire framework is end-to-end, thus allowing for efficient training and inference. $\hat{M}_{i}$ denotes the $i$-th predicted motion and $M_{i}$ denotes the $i$-th reference motion.
@@
In contrast to rule-based methods, deep learning-based approaches utilize data as the primary source to generate 3D face animations. At present, the majority of 3D face animation generation methods are based on speech-driven techniques. In an early work, VOCA (Cudeiro et al. 2019) proposes a temporal convolutional neural network that takes speech and a silent 3D mesh template as input to generate realistic 3D face animations. It also provides an open-source 4D dataset, VOCASET (Cudeiro et al. 2019), as the data base for subsequent studies. MeshTalk (Richard et al. 2021), FaceFormer (Fan et al. 2022), and CodeTalker (Xing et al. 2023) all introduce motion priors. FaceFormer focuses on the long temporal sequences and successfully uses a transformer decoder (Vaswani et al. 2017) to obtain contextual information to generate sequential mesh sequences. CodeTalker adopts VQ-VAE (Van Den Oord, Vinyals et al. 2017) to learn discrete motion priors for generating 3D talking faces for overly smooth facial motions. MeshTalk notes the importance of facial motion in non-lip regions and focuses more on the upper half of the face than VOCA. It adopts categorical latent space to learn discrete motion priors. Wu et al. (2023) goes a step further by focusing on composite and regional facial movements. SelfTalk (Peng et al. 2023a) introduces consistency loss to improve the generation quality.
@@
All these approaches, however, have mainly focused on lip synchronisation but have not paid attention to the correlation between emotions and facial expressions. To address this problem, EmoTalk (Peng et al. 2023b) proposes an end-to-end speech-driven emotional content disentanglement model. However, the transformer part of EmoTalk only uses emotion as the driving source without content. EMOTE (Daněček et al. 2023) also presents an approach of speech-driven content-emotion disentanglement but the effectiveness is limited by the FLAME parameters. More recently, advanced techniques (Stan, Haque, and Yumak 2023; Chen et al. 2023; Sun et al. 2023) introduce diffusion model from 2D image generation into 3D facial animation.
Spiral Convolution
@@
Defferrard et al. (2016) introduces a graph convolutional network (GCN) based on spectral filtering for non-Euclidean graph structure with the same linear complexity as the classical CNN. Kolotouros et al. (2020) devises a simple model consisting of an encoder and a decoder with spiral convolution added to the decoder, which can directly learn the mapping relationships from 2D images to 3D meshes. Although the graph convolution operator is in principle more fundamental and theoretical, it is not intuitive enough for dealing with 3D mesh vertices in the spatial domain. Therefore, Masci et al. (2015) proposes spatial domain graph convolution with sampling of the graph signal. Lim et al. (2018) proposes spiral convolution based on graph convolution to handle 3D mesh vertices in the spatial domain. Gong et al. (2019) introduces SpiralNet++, which utilizes truncated spiral lines to constrain the number of sampled vertices, while incorporating hollow spiral lines to enhance the receptive field. Inspired by these previous works, we use spiral convolution in a 3D facial animation generation task to allow EmoFace to better learn the spatial as well as temporal associations of mesh vertices.
Methods
Overview
@@ Methods
In order to generate vivid emotional 3D talking faces, we propose EmoFace, a model that can sufficiently learn emotions from speech. The input of EmoFace consists of the speech sequence $A_{1:T}=(a_{1},a_{2},{\ldots},a_{T})$, the emotional level $l\in\mathbb{R}^{2}$, the speaking style $s\in\mathbb{R}^{n}$ and the character template $Y_{c}\in\mathbb{R}^{V\times 3}$, where $n$ is the number of speaking styles, $T$ is the length of the mesh sequence, and $V$ is the number of mesh vertices. The emotion level and the speaking style are encoded as one-hot vectors. Our model contains two branches to disentangle and extract the emotion-related features $F_{e}$ and content-related features $F_{c}$ from the speech sequence $A_{1:T}$, respectively, driving the given template vertices to generate offsets $\Delta M^{e}_{1:T}=(\Delta m^{e}_{1},\Delta m^{e}_{2},{\ldots},\Delta m^{e}_{T})$ and $\Delta M^{c}_{1:T}=(\Delta m^{c}_{1},\Delta m^{c}_{2},{\ldots},\Delta m^{c}_{T})$. Finally the output $\Delta M_{1:T}$ is obtained by fusing the predictions of the two branches using the Mesh Attention module. The complete pipeline is shown in Figure 2.
Emotion-Content Disentanglement Algorithm
Inspired by Ji et al. (2021) and Peng et al. (2023b), we use a similar structure to disentangle emotion-related and content-related features to reduce the difficulty of learning the mapping between speech and emotional facial expressions. Firstly, our model consists of two branches that use two pre-trained speech feature extractors wav2vec2.0 (Baevski et al. 2020) as encoders to extract emotional features and content features respectively, where temporal convolutional network (TCN) layer is fixed during fine-tuning. Then, our emotion and content branches use a transformer decoder with the same structure of FaceFormer, which contains a cross-attention module to align the audio and motion modalities as well as a self-attention module to learn the dependencies between each frame of the past facial motions, and improve the ability to generalise to longer sequences. The difference from previous work is that we expect both the emotion branch and the content branch to be as strong as possible in modeling the speech-mesh mapping relationship, so our emotion branch and content branch make separate predictions for the mesh offsets of the next frame and supervised by the groundtruth. However, both branches have significant limitations in their respective abilities to model facial motions. On the one hand, although the features extracted from the content branch are strongly correlated with the lips, the inability to extract long-term emotional features and the lack of the type and intensity of emotion are not conducive to predicting the magnitude of mouth opening. On the other hand, while the features extracted from the emotion branch are strongly correlated with the motions of the eyes and the face expressions, the lack of short-term content features results in insufficiently accurate predictions of the lips. Therefore, we use Mesh Attention to fuse the results of the two branches in temporal and spatial domains to obtain the final prediction.
In the training process, we set a pseudo-training pair of samples with the same content but different emotions. The emotion features and content features are extracted respectively and the two emotion features are exchanged as input of transformer decoder to implement cross reconstruction, as shown in Figure 3. By separating and re-coupling the emotional and content features, the long-term emotional information and short-term content information of speech are forced to be decoupled. The entire training process is described in detail as follows:
$F_{c1}=E_{c}(A_{c1,e1}),F_{e1}=E_{e}(A_{c1,e1}),F_{e2}=E_{e}(A_{c1,e2}),$
$M^{c}_{c1,e1}=M^{c}_{c1,e2}=D_{c}(F_{c1}),$
$M^{e}_{c1,e1}=D_{e}(F_{e1}),M^{e}_{c1,e2}=D_{e}(F_{e2}),$
$M_{c1,e1}=\epsilon(M^{e}_{c1,e1},M^{c}_{c1,e1}),M_{c1,e2}=\epsilon(M^{e}_{c1,e2},M^{c}_{c1,e2}),$ (1)
$\epsilon$ is the Mesh Attention module, $A_{cx,ey}$ represents audio with content $x$ and emotion $y$, $E$ and $D$ are the encoder and decoder respectively, and $M$ represents the Mesh sequence. Our training process is end-to-end, enabling the model with better constraints.
Figure 3: Emotion and content Disentangling Encoder. Various inputs of speech, conveying same content and different emotions, are processed to generate cross-reconstructed mesh vertex offsets representing distinct combinations of facial expressions. Supervisions are added to both two branches and the final output.
Mesh Attention

### v3

# EmoFace: Emotion-Content Disentangled Speech-Driven 3D Talking Face Animation

@@ Abstract
The creation of increasingly vivid 3D talking face has become a hot topic in recent years. Currently, most speech-driven works focus on lip synchronisation but neglect to effectively capture the correlations between emotions and facial motions. To address this problem, we propose a two-stream network called EmoFace, which consists of an emotion branch and a content branch. EmoFace employs a novel Mesh Attention mechanism to analyse and fuse the emotion features and content features. Particularly, a newly designed spatio-temporal graph-based convolution, SpiralConv3D, is used in Mesh Attention to learn potential temporal and spatial feature dependencies between mesh vertices. In addition, to the best of our knowledge, it is the first time to introduce a new self-growing training scheme with intermediate supervision to dynamically adjust the ratio of groundtruth adopted in the 3D face animation task. Comprehensive quantitative and qualitative evaluations on our high-quality 3D emotional facial animation dataset, 3D-RAVDESS ($4.8863\times 10^{-5}$mm for LVE and $0.9509\times 10^{-5}$mm for EVE), together with the public dataset VOCASET ($2.8669\times 10^{-5}$mm for LVE and $0.4664\times 10^{-5}$mm for EVE), demonstrate that our approach achieves state-of-the-art performance.
EmoFace: Emotion-Content Disentangled Speech-Driven
3D Talking Face Animation
Yihong Lin1*, Liang Peng2*, Zhaoxin Fan3,4, Xianjia Wu2, Jianqiao Hu1, Xiandong Li2$\dagger$, Wenxiong Kang1$\dagger$, Songju Lei5, 1South China University of Technology, 2Huawei Cloud, 3Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing, School of Artificial Intelligence, Beihang University, 4Hangzhou International Innovation Institute, Beihang University, 5Nanjing University lxdphys@smail.nju.edu.cn auwxkang@scut.edu.cn
1 Introduction
Generating realistic 2D/3D face animations has received great attention in a number of fields, including film production, computer games, virtual reality, education, etc. Ping et al. (2013); Edwards et al. (2016); Wohlgenannt et al. (2020). Recent methods based on deep learning Karras et al. (2017); Cudeiro et al. (2019); Richard et al. (2021); Fan et al. (2022); Xing et al. (2023); Peng et al. (2023a); Wu et al. (2023); Stan et al. (2023) produce impressive 3D face animations with significant savings in time and labour compared to manual production, making them preferred for academic research and commercial exploration. However, there are still some situations remained to be adequately addressed, for example, the relationship between emotions and facial expressions. Most of the preceding speech-driven methods Cudeiro et al. (2019); Richard et al. (2021); Fan et al. (2022); Xing et al. (2023); Peng et al. (2023a) focus more on achieving high-quality lip synchronisation, while ignoring the impact of emotion on facial animations. Recognizing the significance of emotion, EmoTalk Peng et al. (2023b) and EMOTE Daněček et al. (2023) disentangle emotional and content information in speech to generate high-quality emotional 3D facial animations. Nevertheless, EmoTalk only uses emotion as the driving source, regardless of content, thereby limiting its performance. EMOTE doesn’t achieve end-to-end training and adopts a FLAME Li et al. (2017) parameter-based method that theoretically does not control motion as accurately as vertex-based methods because the abstract coefficient estimation ignores spatial correlations.
@@ 1 Introduction
In this paper, we propose a novel vertex-based autoregressive emotional speech-driven 3D face animation approach. We first disentangle emotion and content from speech and extract their features separately, then a latent space decoder is used to obtain the emotion-based and content-based vertex offsets, which are eventually integrated by Mesh Attention to predict the final offsets. Due to that the fusion weights for the emotion branch and the content branch are variable in both temporal and spatial domains, we add a 3D graph-based convolution operator SpiralConv3D in Mesh Attention for effective feature exaction. Moreover, during the training process, it is observed that the autoregression scheme of the transformer decoder results in error accumulation, which ultimately leads to a deterioration in the output and increases the training time. Although, existing teacher-forcing scheme can help to alleviate this problem, exposure bias may occur when switching from training to inference due to inability to access real historical data. Inspired by scheduled sampling Mihaylova and Martins (2019), we propose a self-growing scheme that gradually adjusts the ratio of groundtruth provided during training. In order to better train our emotion-content disentanglement model, we construct a high-quality dataset 3D-RAVDESS by reconstructing reliable 3D faces from 2D dataset RAVDESS Livingstone and Russo (2018). Extensive qualitative and quantitative experiments demonstrate that our method outperforms current state-of-the-art methods for better generation of facial expressions and lip synchronisation. In summary, our main contributions are as follows:
@@
  • We provide a new high-quality emotional speech-driven dataset, 3D-RAVDESS.
@@
  • We design a two-stream network, EmoFace, to obtain the emotion-based and content-based vertex offsets, respectively, which are dynamically fused by a novel Mesh Attention with SpiralConv3D that efficiently extracts features in the spatio-temporal domain.
  • We introduce intermediate supervision to guide two branches respectively and propose a self-growing training scheme to gradually increase the difficulty of prediction, thus improving the robustness of the model.
  • Extensive experiments demonstrate the superiority of EmoFace over existing SOTA methods in terms of full face realism, emotion expression and lip synchronisation on both 3D-RAVDESS and VOCASET.
2 Related Work
2.1 Speech-Driven 3D Talking Face
@@
Speech-driven 3D talking face generation is a task to generate realistic facial animations based on speech Karras et al. (2017); Lahiri et al. (2021); Pham et al. (2017); Taylor et al. (2017). A number of rule-based approaches have been employed to generate 3D motions, including the capture of the relationship between visemes Mattheyses and Verhelst (2015) and facial action units (FAUs) Ekman and Friesen (1978), as well as the establishment of a phoneme-viseme mapping that has yielded highly promising results in early studies Edwards et al. (2016); Xu et al. (2013). Nevertheless, these methods do not easily transfer to new faces as they heavily rely on manual crafting and consume an excessive amount of time.
@@
In contrast to rule-based methods, deep learning-based 3D face animation methods resort to a data-driven framework. Most of them take both speech and a static 3D mesh template as input to generate realistic 3D face animations. In an early work, VOCA Cudeiro et al. (2019) proposes a temporal convolutional neural network with an open-source 4D dataset VOCASET, Cudeiro et al. (2019), which has become a valuable resource for subsequent studies. Building on this foundation, MeshTalk Richard et al. (2021), FaceFormer Fan et al. (2022), and CodeTalker Xing et al. (2023) incorporate motion priors to improve animation quality. MeshTalk notes the importance of facial motion in audio-uncorrelated regions and adopts categorical latent space to learn discrete motion priors. FaceFormer focuses on the long temporal sequences and successfully uses a transformer decoder Vaswani et al. (2017) to obtain contextual information to generate sequential mesh sequences. CodeTalker introduces VQ-VAE Van Den Oord et al. (2017) to learn discrete motion priors for smooth facial motion generation. More recently, SelfTalk Peng et al. (2023a) proposes consistency loss to improve the generation quality. Other works Stan et al. (2023); Chen et al. (2023); Sun et al. (2024) introduce diffusion model into 3D facial animation to enhance motion diversity.
@@
However, all these approaches have mainly focused on lip synchronisation but have not paid attention to the correlation between emotions and facial expressions. To address this problem, EmoTalk Peng et al. (2023b) and EMOTE Daněček et al. (2023) both propose speech-driven content-emotion disentanglement pipelines. Nevertheless, EmoTalk does not utilize content information and only focuses on the driving effect of emotional information, while the effectiveness and efficiency of EMOTE are limited by FLAME parameters and two-stage training pipelines, respectively.
Figure 1: The overall framework of EmoFace. The emotion and content branches disentangle the information in speech, while Mesh Attention fuses these two branches to obtain the final result. The entire framework is end-to-end, thus allowing for efficient training and inference. $\hat{M}_{i}$ denotes the $i$-th predicted motion and $M_{i}$ denotes the $i$-th reference motion. $k$ and $\delta$ represent the spatial and the temporal neighbourhood of SpiralConv3D respectively.
2.2 Spiral Convolution
@@
Defferrard et al. Defferrard et al. (2016) present a graph convolutional network (GCN) based on spectral filtering for non-Euclidean graph structure with the same linear complexity as the classical CNN. Kolotouros et al. Kulon et al. (2020) devise a simple model consisting of an encoder and a decoder with spiral convolution added to the decoder, which can directly learn the mapping relationships from 2D images to 3D meshes. Since the graph convolution operator is intuitive enough for dealing with 3D mesh vertices in the spatial domain, Masci et al. Masci et al. (2015) advance the idea of spatial graph convolution with sampling of the graph signal. Lim et al. Lim et al. (2018) propose spiral convolution based on graph convolution to handle 3D mesh vertices in the spatial domain. Gong et al. Gong et al. (2019) introduce SpiralNet++, which utilizes truncated spiral lines to constrain the number of sampled vertices, while incorporating hollow spiral lines to enhance the receptive field. Inspired by these previous works, we use spiral convolution in a 3D facial animation generation task to allow EmoFace to better learn the spatial as well as temporal associations of mesh vertices.
3 Methods
3.1 Overview
@@
The overall pipeline is shown in Figure 1. In order to generate vivid emotional 3D talking faces, we propose EmoFace, a model that can sufficiently learn emotions from speech and generate talking face with rich expressions. The input of EmoFace consists of the speech sequence $A_{1:T}=(a_{1},a_{2},{\ldots},a_{T})$, the emotional level $l\in\mathbb{R}^{2}$, the speaking style $s\in\mathbb{R}^{n}$ and the character template $Y_{c}\in\mathbb{R}^{V\times 3}$, where $n$ is the number of speaking styles, $T$ is the length of the mesh sequence, and $V$ is the number of mesh vertices. The emotion level and the speaking style are encoded as one-hot vectors. Inspired by Ji et al. (2021) and Peng et al. (2023b), we use a similar emotion-content disentanglement module that uses pre-trained speech feature extractors wav2vec2.0 Baevski et al. (2020) to reduce the difficulty of learning the mapping between speech and emotional facial expressions. Specifically, the emotion branch and the content branch firstly disentangle emotion-related features $F_{e}$ and content-related features $F_{c}$ from the speech sequence $A_{1:T}$. Then, emotion-related features $F_{e}$ and content-related features $F_{c}$ drive the given template vertices to generate offsets $\Delta M^{e}_{1:T}=(\Delta m^{e}_{1},\Delta m^{e}_{2},{\ldots},\Delta m^{e}_{T})$ and $\Delta M^{c}_{1:T}=(\Delta m^{c}_{1},\Delta m^{c}_{2},{\ldots},\Delta m^{c}_{T})$, respectively, using the same transformer decoder as FaceFormer, which contains a multi-head cross-attention module to align the audio and motion modalities as well as a multi-head self-attention module to learn the dependencies between each frame of the past facial motions. Finally, Mesh Attention is provided to fuse the predictions of the two branches in temporal and spatial domains to obtain the final prediction $\Delta M_{1:T}$.
3.2 Mesh Attention
In order to fuse the prediction results of the emotion branch and content branch, we propose a novel spatio-temporal attention module Mesh Attention that analyses the fusion weight. Different from conventional MLP-based motion decoding methods that suffer from mesh Non-Euclidean topology destruction and explicit geometric constraint absence, Mesh Attention employs the core operator SpiralConv3D to directly model spatial vertex connectivity within frames and temporal vertex correspondence across frames. This approach performs spatiotemporal aggregation of neighboring vertices’ geometric and semantic features while maintaining the mesh’s local manifold structure and physical constraints (e.g., smoothness, symmetry), ultimately generating a Temporal Mesh Attention Map that quantifies the impact of two branches on final predictions.